### v1

@@ 6 Conclusion
In this paper, we study the pay and non-pay content of job ads. Our first set of results is derived from analyzing the texts of a comprehensive database of job ads. We find that the texts of these ads specify many job attributes relevant to workers’ preferences over alternative employers. Almost every job ad mentions at least one of these attributes with an average of 6.6 attributes per ad. By considering a wide range of advertised job attributes, our analysis adds to prior work focusing on specific content, such as salary information (Batra et al., 2023), remote work (Hansen et al., 2023), or flexible working arrangements (Adams-Prassl et al., 2023a).
We also recover workers’ valuation for alternative employers using a structural model estimated on matched employer-employee data. Our second set of results follows from linking these estimates to the pay and non-pay content of job ads. High-pay employers are more likely to advertise aspects of the compensation package, while high-amenity employers more often mention working arrangements, such as contract duration and hours worked, in their postings. We confirm using several alternative metrics that the content of job ads has predictive power for workers’ valuation of employers above other observable characteristics, such as industries and occupations. Taken together, these results suggest that the pay and non-pay content of job vacancies represents one reliable source of information for workers assessing their options in the labor market.
@@
Finally, our results have implications for the design of policies related to the information contained in job ads. Several jurisdictions have passed legislation mandating employers to disclose a salary range in their job openings. Our analysis suggests that there is scope to regulate the provision of additional job characteristics besides pay, such as information on working arrangements. Similarly, these results can be useful for the design of online job platforms in that they provide guidance on the attributes employers should be encouraged to provide in their postings.
@@ Appendix A Additional Tables and Figures
Table A.1: Overview of Sample Selection: Five-Year Periods, Cluster-Level Match.
@@
Panel A: 2000-2004
All Observations $339,791$ (100%) $210,958$ (100%) $2,468,402$ (100%) $9,820,699$ (100%)
Observe Vacancy Text $330,374$ (97.2%) $87,790$ (41.6%) $2,077,549$ (84.2%) $8,488,411$ (86.4%)
Excluding Staffing Agencies $323,332$ (95.2%) $87,789$ (41.6%) $2,077,548$ (84.2%) $8,488,320$ (86.4%)
Strongly Connected Set $221,999$ (65.3%) $54,759$ (26.0%) $1,934,829$ (78.4%) $7,915,843$ (80.6%)
All Sample Restrictions $209,798$ (61.7%) $54,752$ (26.0%) $1,934,476$ (78.4%) $7,914,290$ (80.6%)
Panel B: 2005-2009
All Observations $948,589$ (100%) $228,287$ (100%) $2,663,696$ (100%) $10,706,270$ (100%)
Observe Vacancy Text $917,501$ (96.7%) $119,000$ (52.1%) $2,389,467$ (89.7%) $9,727,646$ (90.9%)
Excluding Staffing Agencies $850,809$ (89.7%) $119,000$ (52.1%) $2,389,467$ (89.7%) $9,727,646$ (90.9%)
Strongly Connected Set $708,646$ (74.7%) $61,459$ (26.9%) $2,134,368$ (80.1%) $8,802,142$ (82.2%)
All Sample Restrictions $631,441$ (66.6%) $61,453$ (26.9%) $2,134,333$ (80.1%) $8,801,945$ (82.2%)
Panel C: 2010-2014
All Observations $734,996$ (100%) $237,677$ (100%) $2,854,890$ (100%) $11,797,654$ (100%)
Observe Vacancy Text $702,292$ (95.6%) $110,132$ (46.3%) $2,513,019$ (88.0%) $10,570,006$ (89.6%)
Excluding Staffing Agencies $670,108$ (91.2%) $110,132$ (46.3%) $2,513,019$ (88.0%) $10,570,006$ (89.6%)
Strongly Connected Set $572,520$ (77.9%) $62,934$ (26.5%) $2,264,686$ (79.3%) $9,632,981$ (81.7%)
All Sample Restrictions $521,661$ (71.0%) $62,933$ (26.5%) $2,264,662$ (79.3%) $9,632,258$ (81.6%)
Panel D: 2015-2019
All Observations $915,790$ (100%) $301,716$ (100%) $3,123,055$ (100%) $13,278,335$ (100%)
Observe Vacancy Text $863,478$ (94.3%) $127,913$ (42.4%) $2,733,050$ (87.5%) $11,898,501$ (89.6%)
Excluding Staffing Agencies $774,076$ (84.5%) $127,904$ (42.4%) $2,732,297$ (87.5%) $11,895,973$ (89.6%)
Strongly Connected Set $744,108$ (81.3%) $78,133$ (25.9%) $2,500,325$ (80.1%) $10,977,778$ (82.7%)
All Sample Restrictions $624,992$ (68.2%) $78,118$ (25.9%) $2,499,506$ (80.0%) $10,974,896$ (82.7%)
Panel E: 2000-2019
All Observations $2,939,166$ (100%) $472,945$ (100%) 3,867$,23$ 867 , 232 (100%) $45,602,960$ (100%)
Observe Vacancy Text $2,813,645$ (95.7%) $222,739$ (47.1%) $3,722,456$ (96.3%) $40,684,564$ (89.2%)
Excluding Staffing Agencies $2,618,325$ (89.1%) $222,739$ (47.1%) $3,722,316$ (96.3%) $40,681,944$ (89.2%)
Strongly Connected Set $2,247,273$ (76.5%) $126,925$ (26.8%) $3,621,634$ (93.6%) $37,328,744$ (81.9%)
All Sample Restrictions $1,987,892$ (67.6%) $126,914$ (26.8%) $3,621,402$ (93.6%) $37,323,388$ (81.8%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for each of the four five-year periods (Panels A-D) and the combined 2000-2019 period (Panel E). “All Observations” refers to observations with valid establishment identifiers in each five-year period. “Excluding Staffing Agencies” are observations linked to at least one vacancy with observed text not posted by a recruitment or temporary employment agency. “Strongly Connected Set” refers to the strongly connected set of employers used in the structural model. The sample sizes are calculated after matching vacancies to the data on the establishment-cluster level in the corresponding five-year period. In Panel A, vacancy texts are only observed for 2002-2004 due to the availability of data on job ads, and similarly, in Panel E, vacancy texts are observed between 2002 and 2019.
Table A.2: Overview of Sample Selection: Five-Year Periods, Establishment-Level Match.
Job Ads Establishments Workers Worker-Years
(1) (2) (3) (4)
Panel A: 2000-2004
All Observations $339,791$ (100%) $210,957$ (100%) $2,468,402$ (100%) $9,820,699$ (100%)
Observe Vacancy Text $330,374$ (97.2%) $65,578$ (31.1%) $1,613,285$ (65.4%) $6,623,923$ (67.4%)
Excluding Staffing Agencies $323,332$ (95.2%) $65,264$ (30.9%) $1,589,271$ (64.4%) $6,553,017$ (66.7%)
Strongly Connected Set $221,999$ (65.3%) $54,759$ (26.0%) $1,934,829$ (78.4%) $7,915,843$ (80.6%)
All Sample Restrictions $209,798$ (61.7%) $32,342$ (15.3%) $1,446,473$ (58.6%) $5,980,817$ (60.9%)
Panel B: 2005-2009
All Observations $948,589$ (100%) $228,286$ (100%) $2,663,696$ (100%) $10,706,270$ (100%)
Observe Vacancy Text $917,501$ (96.7%) $105,358$ (46.2%) $2,165,790$ (81.3%) $8,822,007$ (82.4%)
Excluding Staffing Agencies $850,809$ (89.7%) $104,434$ (45.7%) $2,123,522$ (79.7%) $8,696,302$ (81.2%)
Strongly Connected Set $708,646$ (74.7%) $61,459$ (26.9%) $2,134,368$ (80.1%) $8,802,142$ (82.2%)
All Sample Restrictions $631,441$ (66.6%) $47,293$ (20.7%) $1,869,473$ (70.2%) $7,774,969$ (72.6%)
Panel C: 2010-2014
All Observations $734,996$ (100%) $237,676$ (100%) $2,854,890$ (100%) $11,797,654$ (100%)
Observe Vacancy Text $702,292$ (95.6%) $92,453$ (38.9%) $2,200,499$ (77.1%) $9,285,465$ (78.7%)
Excluding Staffing Agencies $670,108$ (91.2%) $91,377$ (38.4%) $2,146,382$ (75.2%) $9,113,678$ (77.2%)
Strongly Connected Set $572,520$ (77.9%) $62,934$ (26.5%) $2,264,686$ (79.3%) $9,632,981$ (81.7%)
All Sample Restrictions $521,661$ (71.0%) $44,661$ (18.8%) $1,899,569$ (66.5%) $8,180,234$ (69.3%)
Panel D: 2015-2019
All Observations $915,790$ (100%) $301,715$ (100%) $3,123,055$ (100%) $13,278,335$ (100%)
Observe Vacancy Text $863,478$ (94.3%) $108,730$ (36.0%) $2,456,581$ (78.7%) $10,770,665$ (81.1%)
Excluding Staffing Agencies $774,076$ (84.5%) $107,204$ (35.5%) $2,401,811$ (76.9%) $10,566,082$ (79.6%)
Strongly Connected Set $744,108$ (81.3%) $78,133$ (25.9%) $2,500,325$ (80.1%) $10,977,778$ (82.7%)
All Sample Restrictions $624,992$ (68.2%) $58,153$ (19.3%) $2,170,687$ (69.5%) $9,651,061$ (72.7%)
Panel E: 2000-2019
All Observations $2,939,166$ (100%) $472,945$ (100%) 3,867$,23$ 867 , 232 (100%) $45,602,960$ (100%)
Observe Vacancy Text $2,813,645$ (95.7%) $200,220$ (42.3%) $3,589,923$ (92.8%) $35,502,060$ (77.9%)
Excluding Staffing Agencies $2,618,325$ (89.1%) $197,986$ (41.9%) $3,567,109$ (92.2%) $34,929,080$ (76.6%)
Strongly Connected Set $2,247,273$ (76.5%) $126,925$ (26.8%) $3,621,634$ (93.6%) $37,328,744$ (81.9%)
All Sample Restrictions $1,987,892$ (67.6%) $94,530$ (20.0%) $3,419,517$ (88.4%) $31,587,080$ (69.3%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for each of the four five-year periods (Panels A-D) and the 2000-2019 period (Panel E). “All Observations” refers to observations with valid establishment identifiers in each five-year period. “Excluding Staffing Agencies” are observations linked to at least one vacancy with observed text not posted by a recruitment or temporary employment agency. “Strongly Connected Set” refers to the strongly connected set of employers used in the structural model. The sample sizes in the last row in each panel are calculated after matching vacancies to the data on the establishment level in the corresponding five-year period (i.e., this excludes all establishments without posted vacancies in the corresponding five-year period). In Panel A, vacancy texts are only observed for 2002-2004 due to the availability of data on job ads, and similarly, in Panel E, vacancy texts are observed between 2002 and 2019.
Table A.3: Overview of Establishment Characteristics.
All Establishments A Observe Vacancy Text Excluding Staffing Agencies Strongly Connected Set Both Sample Restrictions
Panel A: Worker Characteristics
Women $43.7$% $49.9$% $49.9$% $51.2$% $51.2$%
College $34.9$% $37.9$% $37.9$% $40.2$% $40.2$%
Hourly Wage $25.34$ $25.16$ $25.15$ $25.45$ $25.45$
Panel B: Establishment Characteristics
Size $13.2$ $22.5$ $22.6$ $30.1$ $30.1$
Public Sector $17.3$% $25.8$% $25.8$% $30.7$% $30.7$%
Panel C: Industry Composition
Agriculture, Forestry and Fishing $2.8$% $1.3$% $1.3$% $1.0$% $1.0$%
@@
Manufacturing $5.8$% $6.4$% $6.5$% $6.6$% $6.6$%
Energy Supply $0.4$% $0.5$% $0.5$% $0.4$% $0.4$%
Water Supply $0.5$% $0.6$% $0.6$% $0.6$% $0.6$%
Construction $10.3$% $8.2$% $8.2$% $8.0$% $8.0$%
Wholesale and Retail Trade $23.3$% $24.6$% $24.6$% $22.9$% $22.9$%
Transportation and Storage $5.5$% $4.5$% $4.5$% $4.6$% $4.6$%
Accommodation and Food Service $4.2$% $4.9$% $4.9$% $4.7$% $4.7$%
Information and Communication $3.1$% $2.8$% $2.8$% $2.9$% $2.9$%
Financial and Insurance $1.4$% $1.5$% $1.5$% $1.5$% $1.5$%
Real Estate $3.4$% $1.5$% $1.5$% $1.1$% $1.1$%
Professional, Scientific and Technical $8.2$% $5.8$% $5.8$% $5.1$% $5.1$%
@@
Public Administration and Defense $2.9$% $4.2$% $4.2$% $4.3$% $4.3$%
Education $3.8$% $5.7$% $5.7$% $7.1$% $7.1$%
Health and Social Work $12.9$% $17.3$% $17.4$% $20.0$% $20.0$%
Arts, Entertainment and Recreation $2.0$% $1.8$% $1.8$% $1.7$% $1.7$%
Other Service Activities $4.3$% $3.6$% $3.6$% $2.5$% $2.5$%
Other Industries $0.3$% $0.0$% $0.0$% $0.0$% $0.0$%
Note: This table documents how the sample restrictions imposed in the estimation affects the composition of establishments between 2000 and 2019. Each panel shows averages over establishment-years, meaning that employers that are present in all years are implicitly weighted more than employer that are present in half of the years. Panel A shows average worker characteristics across establishment-years, weighting workers in small employers more than large employers. The first row, for example, is constructed by first calculating the average share of women within employer-years, and then averaging over employer-years. Panel C documents industry compositions using Classification of Economic Activities in the European Community (NACE).
@@
   –– Competitive Pay 14.9 17.5 19.2 94.2 88.6 80.5
   –– Collective Agreement Pay 23.3 31.5 33.8 87.8 84.1 78.5
   –– Incentive Pay Scheme 2.6 2.0 1.5 98.5 50.0 66.7
@@
   –– Pension Scheme 28.6 34.8 32.8 98.0 94.2 100.0
   –– Insurance Scheme 18.0 23.8 21.5 96.2 87.4 96.5
   –– Mortgage Possibility 1.3 2.0 3.8 98.2 100.0 53.3
@@
   –– Good Career Paths 11.9 16.8 6.2 85.0 23.9 64.0
   –– On-the-Job Training 39.2 46.8 39.2 78.0 68.4 81.5
@@
   –– Full-time Contract 58.8 55.8 53.2 95.0 93.3 97.7
   –– Part-time Contract 30.0 34.0 28.2 91.2 78.7 94.7
   –– Full-time/Part-time Choice 4.0 3.0 3.0 94.0 0.0 0.0
   –– On-call Employment 3.3 3.8 2.2 98.0 53.3 88.9
@@
   –– Possibility to Work Flexible Hours 6.1 6.8 7.5 98.2 92.6 83.3
   –– Regular Daytime Work Schedule 14.2 8.8 8.8 97.0 82.9 82.9
   –– Exempt from Work Hour Regulations 0.4 1.2 0.2 98.5 0.0 0.0
@@
   –– Shift Work 14.7 15.0 11.8 95.8 75.0 95.7
   –– Weekend/Evening/Night Work 17.4 21.8 17.0 94.8 77.0 98.5
@@
   –– Social Environment 39.0 44.0 40.8 80.8 74.4 80.4
   –– Good Colleagues 20.4 26.0 28.2 77.2 60.6 55.8
   –– Possibility for Remote Work 0.3 0.5 0.5 100.0 100.0 100.0
   –– Shared Office Space 0.1 0.0 0.0 100.0
   –– Inclusive Work-life Scheme 6.7 7.2 6.0 98.2 79.3 95.8
@@
   –– Interesting Tasks 32.4 38.2 22.2 82.5 56.2 96.6
   –– Challenging Tasks 26.9 32.0 19.0 85.0 56.2 94.7
   –– Variation in Tasks 10.3 10.2 15.8 87.5 65.9 42.9
   –– Responsibility in Job 1.5 1.8 10.2 90.5 71.4 12.2
   –– Independence in Performing Tasks 2.4 2.8 5.0 95.8 63.6 35.0
   –– Involves Leadership Responsibility 11.1 13.5 9.0 87.0 35.2 52.8
   –– Work Involves Travelling 7.8 10.2 3.5 93.2 34.1 100.0
@@
   –– Company Gym or Sports Team 3.5 5.0 6.5 95.5 70.0 53.8
@@
   –– Any Welfare Scheme 4.0 4.8 2.5 97.2 47.4 90.0
   –– Company Cabin 1.7 2.0 3.2 98.8 100.0 61.5
   –– Occupational Health Service 1.3 1.2 1.2 100.0 100.0 100.0
@@
   –– Flexible/Extended Holidays 0.4 0.0 0.0 100.0
Number of Job Ads 915,790 400 400 400
@@
By Gender: By Education:
Pooled Sample Men Women Non-College College
(1) (2) (3) (4) (5)
@@
Posted Overall Value ($\tilde{V}_{j}$) -0.154∗∗∗ -0.108∗∗∗ -0.136∗∗∗ -0.121∗∗∗ -0.122∗∗∗
(0.027) (0.025) (0.023) (0.017) (0.018)
@@
Posted Pay Premium ($\tilde{\psi}_{j}$) -0.402∗∗ -0.761∗∗∗ -0.363 -0.378 -0.549∗∗
(0.184) (0.211) (0.187) (0.207) (0.268)
Posted Non-Pay Value ($\tilde{a}_{j}$) -0.573∗∗∗ -0.552∗∗∗ -0.722∗∗∗ -0.430∗∗∗ -0.810∗∗∗
(0.159) (0.173) (0.132) (0.145) (0.175)
Posted Job Security Value ($\tilde{s}_{j}$) -0.086∗∗∗ -0.096∗∗∗ -$0∗∗$082∗∗∗ -0.078∗∗∗
(0.018) (0.018) (0.015) (0.015) (0.021)
@@
Employer Clusters 1,559 858 850 937 761
Establishments 78,104 42,921 42,479 46,834 38,103
Ads 624,875 468,306 507,506 478,109 499,423
Worker-Years 10,974,238 4,931,815 4,956,647 4,973,189 4,664,927
@@
Note: This table documents associations between posted employer values and vacancy duration. Vacancy duration is the number of days between the dates when the ad is posted and unlisted, censored between 1 and 90 days to avoid outliers, averaged at the cluster level before taking logs. We compute the “posted” employer value $(\tilde{V}_{j},\tilde{\psi}_{j},\tilde{a}_{j}$, $\tilde{s}_{j}$) associated with each ad as the prediction from regressions of employer values ($V_{j},\psi_{j},a_{j}$, $s_{j}$) on indicators for posted ad attributes. The “posted” employer pay premium $(\tilde{\psi}_{j}$) and non-pay (amenity) value $(\tilde{a}_{j}$) are measured in logs, so the estimated coefficient for each of these can be interpreted as an elasticity, while the overall employer value $(\tilde{V}_{j}$) and job security ($\tilde{s}_{j}$) are scaled by their standard deviation, so the estimated coefficient for each of these represents a standard deviation increase in employer value. Column (1) shows estimates for the “posted” employer values estimated using the pooled sample, while columns (2)-(5) show the corresponding estimates for each population group. The specification in Panel A focuses only on the overall employer value $(\tilde{V}_{j}$), while Panel B includes its components $(\tilde{\psi}_{j},\tilde{a}_{j}$, $\tilde{s}_{j}$) jointly in the same estimation. Estimations are done at the employer-cluster level for the five-period from 2015 to 2019, and are weighted by the number of worker-years in each cluster and population group. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group).
@@
Note: This figure shows pseudo $R^{2}$ [McFadden, 1974] from separate logistic regressions of binary job attributes detected in job ads posted in Norway between 2002 and 2019 [N=2,813,645], on fixed effects denoting unique combinations of 2-digit industries (90 groups), location indicators (10 groups), 2-digit occupations (43 groups), and 200,220 establishments. We start by including industry fixed effects, continue by including industry×$\times$×location fixed effects, and so on. The last set of regressions controls for 563,170 unique combinations. Location indicators group 422 municipalities into ten groups based on the number of workers such that municipalities assigned to the same group have a similar number of workers, with a specific indicator for job postings in Oslo.

### v2

@@ 6 Conclusion
In this paper, we study the pay and amenity content of job ads. Our first set of results is derived from analyzing the texts of a comprehensive database of job ads. We find that the texts of these ads specify many job attributes relevant to workers’ preferences over alternative employers. Almost every job ad mentions at least one of these attributes with an average of six attributes per ad. By considering a wide range of advertised job attributes, our analysis adds to prior work focusing on specific content, such as pay information (Batra et al., 2023), remote work (Hansen et al., 2023), or flexible work arrangements (Adams-Prassl et al., 2023a).
We then recover workers’ valuation for alternative employers using a structural model estimated on matched employer-employee data. Our second set of results follows from linking these estimates to the pay and non-pay content of job ads. High-pay employers are more likely to advertise aspects of the compensation package, while high-amenity employers more often mention working arrangements, such as contract duration and hours worked, in their postings. We confirm using several alternative metrics that the content of job ads has predictive power for workers’ valuation of employers above other observable characteristics, such as industries and occupations. Taken together, these results suggest that the pay and non-pay content of job vacancies represents one reliable source of information for workers assessing their options in the labor market.
@@
Finally, our results have implications for the design of policies related to the information contained in job ads. Several jurisdictions have passed legislation mandating employers to disclose a salary range in their job openings. Our analysis suggests that there is scope to regulate the provision of information on additional job characteristics besides pay, such as working arrangements. Similarly, our results can be useful for the design of online job platforms in that they provide guidance on the attributes employers should be encouraged to provide in their postings.
@@ Appendix A Additional Tables and Figures
Table A.1: Overview of Sample Selection: All Five-Year Periods.
@@
2000-2004
All Observations $339,791$ (100%) $203,084$ (100%) $2,468,397$ (100%) $9,820,695$ (100%)
Text Analysis $326,582$ (96.1%) $79,744$ (39.3%) $2,226,611$ (90.2%) $8,482,931$ (86.4%)
Model Estimation $221,966$ (65.3%) $54,766$ (27.0%) $1,927,755$ (78.1%) $7,271,922$ (74.0%)
Main Analysis $207,747$ (61.1%) $54,488$ (26.8%) $1,915,369$ (77.6%) $7,197,974$ (73.3%)
2005-2009
All Observations $948,589$ (100%) $215,133$ (100%) $2,663,702$ (100%) $10,706,267$ (100%)
Text Analysis $910,636$ (96.0%) $105,908$ (49.2%) $2,483,551$ (93.2%) $9,721,545$ (90.8%)
Model Estimation $708,884$ (74.7%) $61,476$ (28.6%) $2,060,892$ (77.4%) $7,864,068$ (73.5%)
Main Analysis $628,035$ (66.2%) $60,868$ (28.3%) $2,039,348$ (76.6%) $7,744,231$ (72.3%)
2010-2014
All Observations $734,996$ (100%) $228,191$ (100%) $2,854,890$ (100%) $11,797,651$ (100%)
Text Analysis $698,923$ (95.1%) $100,721$ (44.1%) $2,622,071$ (91.8%) $10,564,100$ (89.5%)
Model Estimation $572,180$ (77.8%) $62,925$ (27.6%) $2,157,573$ (75.6%) $8,442,020$ (71.6%)
Main Analysis $519,459$ (70.7%) $62,122$ (27.2%) $2,125,632$ (74.5%) $8,264,304$ (70.1%)
2015-2019
All Observations $915,789$ (100%) $292,327$ (100%) $3,123,052$ (100%) $13,278,337$ (100%)
Text Analysis $860,467$ (94.0%) $118,741$ (40.6%) $2,882,624$ (92.3%) $11,893,847$ (89.6%)
Model Estimation $744,133$ (81.3%) $78,136$ (26.7%) $2,353,355$ (75.4%) $9,436,842$ (71.1%)
Main Analysis $623,341$ (68.1%) $77,164$ (26.4%) $2,324,318$ (74.4%) $9,239,281$ (69.6%)
2000-2019
All Observations $2,939,165$ (100%) $449,570$ (100%) $3,867,230$ (100%) $45,602,952$ (100%)
Text Analysis $2,796,608$ (95.1%) $195,788$ (43.6%) $3,721,622$ (96.2%) $40,662,424$ (89.2%)
Model Estimation $2,247,163$ (76.5%) $126,931$ (28.2%) $3,335,908$ (86.3%) $33,014,852$ (72.4%)
Main Analysis $1,978,582$ (67.3%) $125,504$ (27.9%) $3,312,618$ (85.7%) $32,445,790$ (71.1%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for each of the four five-year periods and the 2000-2019 period. The first row (“All Observations”) in each panel shows the total number of observations in our data on job ads (Column 1) and matched employer-employee data (Columns 2-4) for each five-year period, prior to the various sample selection steps. The remaining three rows in each panel summarize the sample restrictions imposed in the three separate estimation stages. “Text Analysis” refers to the sample of job ads with non-missing text content and at least one section written in Norwegian. “Model Estimation” restricts the sample to the strongly connected set of employers with workers between 20 and 60 years old (inclusive). “Main Analysis” imposes both of the previous sets of restrictions, and additionally removes recruitment and employment agencies. Job ads are only observed between 2002 and 2019 due to data availability, meaning that the number of observations on job ads in the top and bottom panel refers to the number of job ads posted in 2002-2004 and 2002-2019, respectively.
Table A.2: Overview of Sample Composition.
Text Analysis Model Estimation Main Analysis
All Observations A Observe Vacancy Text Prime-age Workers A Strongly Connected Set All Sample Restrictions
(1) (2) (3) (4) (5)
Worker Characteristics:
   Women 43.7% 49.9% 43.7% 51.2% 51.2%
   College 34.9% 37.9% 34.9% 40.2% 40.2%
   Hourly Wage \$25.34 \$25.17 \$25.34 \$25.45 \$25.45
   Number of Workers 3,123,052 2,882,624 2,671,777 2,353,355 2,324,318
   Number of Worker-Years 13,278,337 11,893,847 11,253,721 9,436,842 9,239,281
Establishment Characteristics:
   Size 11.6 19.9 11.6 26.7 26.7
   Public Sector 17.3% 25.8% 17.3% 30.7% 30.7%
   Industry Composition:
    Agriculture, Forestry and Fishing 2.8% 1.3% 2.8% 1.0% 1.0%
@@
    Manufacturing 5.8% 6.5% 5.8% 6.6% 6.6%
    Energy Supply 0.4% 0.5% 0.4% 0.4% 0.4%
    Water Supply 0.5% 0.6% 0.5% 0.6% 0.6%
    Construction 10.3% 8.2% 10.3% 8.0% 8.0%
    Wholesale and Retail Trade 23.3% 24.6% 23.3% 22.9% 22.9%
    Transportation and Storage 5.5% 4.5% 5.5% 4.6% 4.6%
    Accommodation and Food Service 4.2% 4.9% 4.2% 4.7% 4.7%
    Information and Communication 3.1% 2.8% 3.1% 2.9% 2.9%
    Financial and Insurance 1.4% 1.5% 1.4% 1.5% 1.5%
    Real Estate 3.4% 1.5% 3.4% 1.1% 1.1%
    Professional, Scientific and Technical 8.2% 5.8% 8.2% 5.1% 5.1%
@@
    Public Administration and Defence 2.9% 4.2% 2.9% 4.3% 4.3%
    Education 3.8% 5.7% 3.8% 7.1% 7.1%
    Health and Social Work 12.9% 17.4% 12.9% 20.0% 20.0%
    Arts, Entertainment and Recreation 2.0% 1.8% 2.0% 1.7% 1.7%
    Other Service Activities 4.6% 3.6% 4.6% 2.5% 2.5%
   Number of Establishments 292,327 118,741 268,565 78,136 77,164
Note: This table documents how the sample restrictions imposed in the estimation affects the sample composition between 2000 and 2019. The table shows averages over establishment-years, meaning that employers that are present in all years are implicitly weighted more than employer that are present in half of the years. “Worker Characteristics” refers to average worker characteristics across establishment-years, weighting workers in small employers more than large employers. For instance, the first row is constructed by first calculating the average share of women within employer-years, and then averaging over employer-years. The first column refers to the initial sample prior to the various sample selection steps, while the remaining columns refer to the three estimation stages, as in Table 1. Industry composition is based on the Classification of Economic Activities in the European Community (NACE).
@@
   –– Competitive Pay 15.8 17.5 19.2 94.2 88.6 80.5
   –– Collective Agreement Pay 24.3 30.8 33.8 87.0 83.7 76.3
   –– Incentive Pay Scheme 2.7 2.0 1.5 98.5 50.0 66.7
@@
   –– Pension Scheme 30.4 34.8 32.8 98.0 94.2 100.0
   –– Insurance Scheme 19.1 23.8 21.5 96.2 87.4 96.5
   –– Mortgage Possibility 1.4 2.0 3.8 98.2 100.0 53.3
@@
   –– Good Career Paths 12.6 16.8 6.2 85.0 23.9 64.0
   –– On-the-Job Training 41.6 46.8 39.2 78.0 68.4 81.5
@@
   –– Full-time Contract 59.3 55.8 53.2 95.0 93.3 97.7
   –– Part-time Contract 30.9 34.0 28.2 91.2 78.7 94.7
   –– Full-time/Part-time Choice 4.1 3.0 3.0 94.0 0.0 0.0
@@
   –– Possibility to Work Flexible Hours 6.5 6.8 7.5 98.2 92.6 83.3
   –– Regular Daytime Work Schedule 14.0 8.8 8.8 97.0 82.9 82.9
   –– Exempt from Work Hour Regulations 0.5 1.2 0.2 98.5 0.0 0.0
@@
   –– Shift Work 15.4 15.0 11.8 95.8 75.0 95.7
   –– Weekend/Evening/Night Work 18.4 21.8 17.0 94.8 77.0 98.5
   –– On-call Employment 3.4 3.8 2.2 98.0 53.3 88.9
@@
   –– Social Environment 41.4 44.0 40.8 80.8 74.4 80.4
   –– Good Colleagues 21.7 26.0 28.2 77.2 60.6 55.8
   –– Possibility for Remote Work 0.4 0.5 0.5 100.0 100.0 100.0
   –– Shared Office Space 0.2 0.0 0.0 100.0
   –– Inclusive Work-life Scheme 7.1 7.2 6.0 98.2 79.3 95.8
@@
   –– Interesting Tasks 34.5 38.2 22.2 82.5 56.2 96.6
   –– Challenging Tasks 28.6 32.0 19.0 85.0 56.2 94.7
   –– Variation in Tasks 10.9 10.2 15.8 87.5 65.9 42.9
   –– Responsibility in Job 1.6 1.8 10.2 90.5 71.4 12.2
   –– Independence in Performing Tasks 2.5 2.8 5.0 95.8 63.6 35.0
   –– Involves Leadership Responsibility 11.8 13.5 9.0 87.0 35.2 52.8
   –– Work Involves Travelling 8.2 10.2 3.5 93.2 34.1 100.0
@@
   –– Company Gym or Sports Team 3.7 5.0 6.5 95.5 70.0 53.8
@@
   –– Any Welfare Scheme 4.2 4.8 2.5 97.2 47.4 90.0
   –– Company Cabin 1.9 2.0 3.2 98.8 100.0 61.5
   –– Occupational Health Service 1.4 1.2 1.2 100.0 100.0 100.0
@@
   –– Flexible/Extended Holidays 0.5 0.0 0.0 100.0
Number of Job Ads 860,467 400 400 400
@@
Posted Overall Value ($\tilde{V}_{j}$) -0.139∗∗∗
@@
Posted Pay Premium ($\tilde{\psi}_{j}$) -0.689∗∗∗
Posted Non-Pay Value ($\tilde{a}_{j}$) -1.032∗∗∗
Posted Job Security Value ($\tilde{s}_{j}$) -0.059∗∗∗
@@
Employer Clusters 1,559
Establishments 77,164
Ads 623,341
Worker-Years 9,239,060
@@
Note: This table documents associations between posted employer values and vacancy duration. Vacancy duration is the number of days between the dates when the ad is posted and unlisted, censored between 1 and 90 days to avoid outliers, averaged at the cluster level before taking logs. We compute the “posted” employer value $(\tilde{V}_{j},\tilde{\psi}_{j},\tilde{a}_{j}$, $\tilde{s}_{j}$) associated with each ad as the prediction from regressions of employer values ($V_{j},\psi_{j},a_{j}$, $s_{j}$) on indicators for posted ad attributes. The “posted” employer pay premium $(\tilde{\psi}_{j}$) and non-pay (amenity) value $(\tilde{a}_{j}$) are measured in logs, so the estimated coefficient for each of these can be interpreted as an elasticity, while the overall employer value $(\tilde{V}_{j}$) and job security ($\tilde{s}_{j}$) are scaled by their standard deviation, so the estimated coefficient for each of these represents a standard deviation increase in employer value. The specification in Panel A focuses only on the overall employer value $(\tilde{V}_{j}$), while Panel B includes its components $(\tilde{\psi}_{j},\tilde{a}_{j}$, $\tilde{s}_{j}$) jointly in the same estimation. Estimations are done at the employer-cluster level for the five-period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group).
@@
Note: This figure shows pseudo $R^{2}$ [McFadden, 1974] from separate logistic regressions of binary job attributes detected in job ads posted in Norway between 2002 and 2019 [N=2,533,559], on fixed effects denoting unique combinations of 2-digit industries (91 groups), location indicators (10 groups), 2-digit occupations (43 groups), and 126,890 establishments. We start by including industry fixed effects, continue by including industry×$\times$×location fixed effects, and so on. The regression excludes 263,049 ads with a unique combination of establishment, occupation, location, and industry indicators, as these observations are precisely predicted by the last set of regressions by construction. The last set of regressions controls for 297,142 unique combinations. Location indicators group 422 municipalities into ten groups based on the number of workers such that municipalities assigned to the same group have a similar number of workers, with a specific indicator for job postings in Oslo.